A proton (q=1.6x10 -19 C) moves 2.0 x 10 -6 m in the direction of electric field that has a magnitude of 2.0 N/C. What is the change in the electrical potential energy associated with the proton?
1
Expert's answer
2020-06-22T11:08:01-0400

The change in the electrical potential energy of the proton


"\Delta W_p=qEd\\
=1.6\times 10^{-19}\times 2.0\times 2.0\times 10^{-6}=6.4\times 10^{-25}\;\rm J"